OpenStudy (anonymous):

Factor the expression x2 − 4x − 12.

OpenStudy (anonymous):

ANSWERS:

OpenStudy (kl0723):

the general rule to use here is to look at the last term in this case (12) and find two numbers that will add or substract to 4 :)

OpenStudy (anonymous):

x^2 - 4x - 12 = 0 (x - 6)(x + 2) = 0 So it's the second option.

OpenStudy (kl0723):

yes :) 6-4 equals your middle term and they both multiply to 12

OpenStudy (anonymous):

You guys are amazing, thanks so much !

OpenStudy (anonymous):

No problem!